# Partial Diacritization: A Context-Contrastive Inference Approach

### v2

@@ 9.2 Partial Diacritization
Similar to this work, previous research has explored the selective diacritization of Arabic text. \@BBOPcitet\@BAP\@BBNAlnefaie and Azmi (2017)\@BBCP harnessed the output of the MADAMIRA morphological analyzer \@BBOPcite\@BAP\@BBNPasha et al. (2014)\@BBCP and leveraged WordNet to generate word candidates for diacritics. By focusing on resolving word ambiguity through statistical and contextual similarity approaches, this research aimed to enhance diacritization effectiveness. \@BBOPcite\@BAP\@BBNAlqahtani et al. (2019)\@BBCP focused on selective homograph disambiguation, proposing methods to automatically identify and mark a subset of words for diacritic restoration. Evaluation of various strategies for ambiguous word selection revealed promising results in downstream applications, such as neural machine translation, part-of-speech tagging, and semantic textual similarity, demonstrating that partial diacritization effectively strikes a balance between homograph disambiguation and mitigating sparsity effects. \@BBOPcitet\@BAP\@BBNEsmail et al. (2022)\@BBCP employ two neural networks are—one considering the entire sentence and the other considering the text read so far—to predict diacritics. Partial diacritization decisions are made based on disagreements between the two networks, favoring the holistic sentence-based approach.
@@ 10 Discussion
In this work, we have presented a novel approach for partial diacritization using context-contrastive prediction. The motivation behind using partial diacritization stems from a large body of literature on the impact of diacritization coverage on the reading process of Arabic text, and the lack of research done in this area.
@@
A significant motivation for our work is grounded in prior research that underscores the substantial influence of diacritization on the reading process of Arabic text. Multiple studies have consistently shown that skilled readers tend to read extensively diacritized text at a slower pace compared to undiacritized text. This phenomenon, supported by research involving diverse age groups and linguistic backgrounds, has significant implications for reading accuracy and comprehension \@BBOPcitep\@BAP\@BBN(Taha, 2016; Ibrahim, 2013; Abu-Leil et al., 2014; Midhwah and Alhawary, 2020; Roman and Pavard, 1987; Hermena et al., 2015; Hallberg, 2022)\@BBCP. Other studies also suggest that diacritization not only affects reading speed but also has a direct impact on reading accuracy \@BBOPcitep\@BAP\@BBN(Roman and Pavard, 1987; Hermena et al., 2015)\@BBCP. In this work, we complement this large body of research by conducting a behavioral experiment that employed our CCPD approach to partially diacritize news headlines from the WikiNews test set. Our results further show that partially diacritized text not only is easier to read than fully diacritized text, but in some cases can lead to better understanding than no diacritization at all.
@@
In this work, we proposed a battery of indicators to gauge the performance of partial diacritization systems. However, further research is needed to gather high-quality partial diacritization benchmarks, allowing the incorporation of additional metrics or indicators for assessing model performance. Evaluating our system’s performance on other languages that employ diacritical marks in their core orthography would clarify the general applicability of our approach.
@@ 11 Conclusion
In conclusion, our CCPD approach for partial diacritization using context-contrastive prediction not only contributes to the field of diacritization but also has far-reaching implications for optimizing the reading experience of Arabic text. Grounded in previous research on the influence of diacritization on reading—further supported by behavioral experiments conducted in this work—our approach paves the way for advancements that benefit readers of diverse backgrounds and abilities. Since our approach integrates seamlessly with existing Arabic diacritization systems, it can be used by used by models trained on domain-specific text as well. We also propose a battery of performance indicators to measure the competence of partial diacritization systems due to the lack of publicly available partial diacritization benchmarks. Finally, we introduce a new diacritization model, TD2—a Transformer adaptation of the D2 model which offers a different performance profile as measured by our proposed indicators. As Arabic NLP continues to evolve, our approach serves as a promising direction for enhancing diacritization and its impact on text accessibility and comprehension.
@@ A.1 Datasets
In this work, we train on data from the Tashkeela \@BBOPcitep\@BAP\@BBN(Fadel et al., 2019)\@BBCP and Ashaar \@BBOPcitep\@BAP\@BBN(Alyafeai et al., 2023)\@BBCP diacritization corpora. We test on the Tashkeela testset. The Tashkeela corpus has been collected mostly from Islamic classical books \@BBOPcite\@BAP\@BBNZerrouki and Balla (2017)\@BBCP and contains mostly classical Arabic sentences. Ashaar is a corpus of Arabic poetry verses covering poems from different eras. Table 4 shows the number of tokens for the Tashkeela and Ashaar555The reported numbers reflect dataset cleaning to keep only Arabic letters and diacritics. datasets.
@@ Appendix B Demo and Examples
We developed an online Demo on Huggingface666huggingface.co/spaces/bkhmsi/Partial-Arabic-Diacritization which allows choice between Full Diacritization, and Partial Diacritization with Hard and Soft modes. See Figures 4(a) and 4(b). The demo supports the D2/TD2 models (trained on Tashkeela and Tashkeela/Ashaar). Since they focus mostly on classical Arabic text, users are advised not to anticipate optimal performance when applying this model to Modern Standard Arabic (MSA) or informal Arabic dialects.
@@ B.1 More Partial Diacritization Examples
[…]

### v3

@@ 9.2 Partial Diacritization
Similar to this work, previous research has explored the selective diacritization of Arabic text. \@BBOPcitet\@BAP\@BBNAlnefaie and Azmi (2017)\@BBCP harnessed the output of the MADAMIRA morphological analyzer \@BBOPcite\@BAP\@BBNPasha et al. (2014)\@BBCP and leveraged WordNet to generate word candidates for diacritics. This work focused on resolving word ambiguity through statistical and contextual similarity approaches to enhance diacritization effectiveness. \@BBOPcitet\@BAP\@BBNAlqahtani et al. (2019a)\@BBCP focused on selective homograph disambiguation, proposing methods to automatically identify and mark a subset of words for diacritic restoration. Evaluation of various strategies for ambiguous word selection revealed promising results in downstream applications, such as neural machine translation, part-of-speech tagging, and semantic textual similarity, demonstrating that partial diacritization effectively strikes a balance between homograph disambiguation and mitigating sparsity effects. \@BBOPcitet\@BAP\@BBNEsmail et al. (2022)\@BBCP employ two neural networks to predict partial diacritics—one considering the entire sentence and the other considering the text read so far. Partial diacritization decisions are made based on disagreements between the two networks, favoring the prediction conditioned on the whole sentence.
@@ 10 Discussion
In this work, we have presented a novel approach for partial diacritization using context-contrastive prediction. The task is motivated by a large body of literature on the impact of diacritization coverage on the reading process of Arabic text, and the scarcity of research into systems to automate it.
@@
A significant motivation for our work rests upon prior research that into the substantial influence of diacritization on the reading process of Arabic text. Multiple studies have consistently shown that skilled readers tend to read extensively diacritized text at a slower pace than undiacritized text. This phenomenon, supported by research involving diverse age groups and linguistic backgrounds, has significant implications for reading accuracy and comprehension \@BBOPcitep\@BAP\@BBN(Taha, 2016; Ibrahim, 2013; Abu-Leil et al., 2014; Midhwah and Alhawary, 2020; Roman and Pavard, 1987; Hermena et al., 2015; Hallberg, 2022)\@BBCP. Other studies also suggest that diacritization not only affects reading speed but also directly impacts reading accuracy \@BBOPcitep\@BAP\@BBN(Roman and Pavard, 1987; Hermena et al., 2015)\@BBCP. In this work, we complement this large body of research by conducting a behavioral experiment that utilized our approach to partially diacritize news headlines from the WikiNews test set. Our results show that partially diacritized text is easier to read than fully diacritized text, and in some cases can lead to better understanding than no diacritization at all.
@@
Further research is needed to gather high-quality benchmark data for partial diacritization, to enable more traditional and direct performance metrics alongside the indicators we propose. One promising direction is evaluating the method on other languages which utilize diacritical marks in their orthographies.
@@ 11 Conclusion
In conclusion, our approach for partial diacritization using context-contrastive prediction contributes to the field of diacritization in an area with far-reaching implications for enhancing the reading experience of Arabic text. Grounded in previous research on the influence of diacritization on reading and further supported by behavioral experiments conducted in this work—our approach paves the way for advancements that benefit readers of diverse backgrounds and abilities. Since our approach integrates seamlessly with existing Arabic diacritization systems, it can be used by models trained on domain-specific text as well. We also propose a battery of performance indicators to gauge the competence of partial diacritization systems using fully-diacritized test data to mitigate the lack of publicly available benchmarks. Finally, we introduce TD2—a Transformer adaptation of the D2 model which offers a different performance profile as shown by our proposed indicators. As Arabic NLP continues to evolve, our approach serves as a promising direction for enhancing diacritization and its impact on text accessibility and comprehension.
Limitations
The method presented in this work is tested on a single human language. While we believe it should be able to generalize to other languages that need similar diacritics restoration, its utility needs further observation. In particular, it may be the case that partial diacritization patterns in languages besides Arabic are different in such a way that the premise of this work no longer holds (wherein ease of guessing is regarded as a major factor in omitting a diacritical mark).
In addition, this method is deliberately kept simple, which builds on an over-simplified view of the human task. Readers guess diacritics when reading, but they may do so while incorporating context via some simple or fast mechanisms. Whether that is at a deep level equivalent to guessing the reading of a single word in isolation remains to be seen. Our method uses non-contextual application, , as a proxy for this process, which is likely relatively close in performance to the same mechanism in a human. Nevertheless, we make no claim that this is indeed the natural mechanism, nor that the proxy exhibits identical performance distribution to the natural mechanism.
@@ A.1 Datasets
In this work, we train on data from the Tashkeela \@BBOPcitep\@BAP\@BBN(Fadel et al., 2019)\@BBCP and Ashaar \@BBOPcitep\@BAP\@BBN(Alyafeai et al., 2023)\@BBCP corpora. We test on the Tashkeela testset. The Tashkeela corpus has been collected mostly from Islamic classical books \@BBOPcite\@BAP\@BBNZerrouki and Balla (2017)\@BBCP and contains mostly classical Arabic sentences. Ashaar is a corpus of Arabic poetry verses covering poems from different eras. Table 4 shows the number of tokens for the Tashkeela and Ashaar666The reported numbers reflect dataset cleaning to keep only Arabic letters and diacritics. datasets.
@@ Appendix B Demo and Examples
We developed an online Demo on Huggingface777huggingface.co/spaces/bkhmsi/Partial-Arabic-Diacritization which allows choice between Full Diacritization, and Partial Diacritization with Hard and Soft modes. See Figures 5(a) and 5(b). The demo supports the D2/TD2 models (trained on Tashkeela and Tashkeela/Ashaar). Since they focus mostly on classical Arabic text, users are advised not to anticipate optimal performance when applying this model to Modern Standard Arabic (MSA) or informal Arabic dialects.
@@ B.1 More Partial Diacritization Examples
The following examples are lines taken from a poem by Rami Mohamed (2016) called “We Will Stay Here”. We applied the models as indicated in the System column. The input text included no diacritics at all. See Table 5.